OpenStudy (anonymous):

Use the distributive property to simplify the expression below: The quantity of negative 2 times x minus 8 all divided by negative 2 x + 4 x − 2 −x + 2 x − 8

OpenStudy (anonymous):

Leon's age can be described as 2 less than 3 times his cousin’s age. If his cousin’s age is represented by the variable "m", write the algebraic expression representing Leon's age using "m".

OpenStudy (anonymous):

(4x - 10)/-2 = 4x/-2 - 10/-2 = -2x + 5

OpenStudy (anonymous):

understood?

OpenStudy (debbieg):

What problem are you doing @kathy0514 ? He has two questions posted in this thread and that isn't either of them. lol

OpenStudy (anonymous):

@Niall

OpenStudy (anonymous):

Yeah, but these are tha answers x + 4 x − 2 −x + 2 x − 8

OpenStudy (debbieg):

Oh - I think she is giving you an example that is similar. Can you follow what she did?

OpenStudy (debbieg):

It's pretty much the same as the one I just helped you with on the other post.... http://openstudy.com/study#/updates/523c8400e4b0fbf3cc7b9d02

OpenStudy (anonymous):

Ohh I see...

OpenStudy (anonymous):

I think I have that one down. Can you guys please help me with this one real quick. Its really important... Leon's age can be described as 2 less than 3 times his cousin’s age. If his cousin’s age is represented by the variable "m", write the algebraic expression representing Leon's age using "m".

OpenStudy (anonymous):

I have about 2 mins to do it

OpenStudy (debbieg):

your set-up is just: \(\Large \dfrac{ -2x-8 }{- 2 }=\dfrac{ -1 }{ 2 }\left(-2x-8\right)\)

OpenStudy (debbieg):

His cousins age is m. What is 3 times his cousin's age?

OpenStudy (anonymous):

3m

OpenStudy (debbieg):

Then, what is 2 less than THAT?

OpenStudy (debbieg):

right - now what is 2 less than that?

OpenStudy (anonymous):

1?

OpenStudy (debbieg):

noooo.... 2 less than (3m) they are not like terms.

OpenStudy (anonymous):

would it be 2 - 3m written out?

OpenStudy (debbieg):

almost - that's closer.... but THAT is 3m less than 2. You want 2 less than 3m.

OpenStudy (debbieg):

If said "2 less than 5", you would say "5-2" right? So what is "2 less than 3m"?

OpenStudy (anonymous):

3m - 2?

OpenStudy (debbieg):

Exactly!

OpenStudy (anonymous):

I see thank you!

OpenStudy (debbieg):

youre welcome :0

OpenStudy (anonymous):

@Niall : Got it ?:)